The following real numbers have decimal expansions as given below. In each case, decide whether they are rational or not. If they are rational, and of the form, p q what can you say about the prime factors of q? 
43.123456789 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

43.123456789
Since it has a terminating decimal expansion, it is a rational number in the form of p/q and q has factors of 2and 5 only.

Thanks
Ram Kushwah
3 years ago
43.123456789 is having terminating decimal expansion
so the number is a rational number
 
Now 43.123456789=43123456789/1000000000
=43123456789/10^9
=43123456789/ 2^9 *5^9
so p=43123456789
 
and q=2^9 *5^9
 
Thus q is in the form of 2^m *5^n
here m=9 and n=9